OpenStudy (emilyjones284):

help! I give medals (:

OpenStudy (emilyjones284):

\[ \sqrt{2bc} \over \sqrt{3}\]

OpenStudy (emilyjones284):

rationalize the denominator and simplify

OpenStudy (anonymous):

well you cant have sqrt on the bottom so we have to cancel them by multiplying\[\frac{ \sqrt{2bc} }{ \sqrt{3} } * \frac{ \sqrt{3} }{ \sqrt{3} } = \frac{ \sqrt{2bc}*\sqrt{3} }{ 3 }\] i hope this help you get started :)

OpenStudy (emilyjones284):

thanks! i got that part but i don't know how to multiply the top part together...thats my problem

OpenStudy (anonymous):

sorry for the wait i was away but if we look at the what we have right now\[\frac{ \sqrt{2bc}*\sqrt{3}}{ 3 }\] only step that i see to take is just multiply the \[\sqrt{2bc}*\sqrt{3}=\sqrt{6bc}\] now our whole looks like \[\frac{ \sqrt{6bc} }{ 3 }\] correct me if im wrong but i dont think you can go any further with this problem since we cant break down 6 into that would be a perfect squared i think we are done

OpenStudy (emilyjones284):

ohhh okay thats all I needed.. thanks a lot! (:

OpenStudy (anonymous):

no problem, i did the best i could :) sorry i could not be more sure :O

OpenStudy (emilyjones284):

its okay! (: